11: What is the length of the edge of a cube with surface area 36 cm2? (BLANK) cm

Answer by dolly(163)

Surface area of a cube = 36 sq cm
AS per formula, the surface area = 6a^2 where 'a' is the edge.
==> 6a^2 = 36
==> a^2 = 36/6
==> a^2 = 6
==> a = sqrt(6)
==> a = 2.45
The edge of the cube = 2.45 cm